Divide 50 into two parts such that the sum of their reciprocals is 1/12. the numbers are
(a)(24,26) (b) (28,22) (c) (27,23) (d) (20,30)

Answer by mananth(16949)

Divide 50 into two parts such that the sum of their reciprocals is 1/12. the numbers are
(a)(24,26) (b) (28,22) (c) (27,23) (d) (20,30)
let one number be x
the other number will be 50-x
1/x + 1/(50-x)= 1/12
50-x+x = x(50-x)/12
600 = 50x-x^2
x^2-50x+600=0
x^2-30x-20x+600=0
x(x-30)-20(x-30)=0
(x-30)(x-20)=0
x= 20or 30 thenumbers are 20 and 30